OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (imstuck):

I don't understand where you got the "+8" part in your first line of the solution up there. Can you explain that to me? I can help you.

OpenStudy (anonymous):

I don't know the whole problem is giving me trouble

OpenStudy (imstuck):

Ok, first of all do you know the rule about the slopes of lines that are parallel to one another?

OpenStudy (anonymous):

I think so

OpenStudy (imstuck):

Ok, what is the rule?

OpenStudy (anonymous):

Sorry i don't no.

OpenStudy (imstuck):

That's ok! The slopes of parallel lines are the same. The rule for perpendicular lines is that their slopes are negative reciprocals. So if the slope for a line is 3, a parallel line will also have the slope of 3. If a line has a slope of 3, a line that is perpendicular to it will have a slope of -1/3.

OpenStudy (imstuck):

So your question asks for the equation of a line that is parallel to the given line. What does that tell you about the slopes of the lines? are they the same or different from each other?

OpenStudy (anonymous):

Different

OpenStudy (imstuck):

Read what I wrote a minute ago again. The slopes of parallel lines are the same. Keep that in mind. That is probably the part that is throwing you off. Let's go on now.

OpenStudy (imstuck):

In order to find the slope of the first line, you have to put it into y=mx+b form. The given equation is not in that form. You have to solve it so y is alone on one side of the equation and the other stuff is on the other side. Can you solve that given equation for y and tell me what you get?

OpenStudy (anonymous):

I don't really get it but i tried and got 19

OpenStudy (imstuck):

That's ok. We will go step by step with the given equation. Just follow this and stop me if you get lost somewhere along the way.

OpenStudy (imstuck):

4x - 3y = -12 is the given equation. -3y = -12 - 4x when you isolate the y term Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Cool. Let's move forward now then! -3y = -12 - 4x is what we have now Because all the terms are negative, and we absolutely HATE negative numbers, we are going to multiply the whole entire mess by -1 to make all the terms positive. Watch:

OpenStudy (imstuck):

-1 (-3y = - 12 - 4x): -1 times -3y is +3y; -1 times -12 is +12; -1 times -4x is +4x. Right? Cuz negative times a negative is a positive.

OpenStudy (imstuck):

So we now have: 3y = 12 + 4x, or to make it closer to the y = mx + b format: 3y = 4x + 12. See that now?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Ok, so the next thing to do is to solve for y. In other words, lose the 3 in front of it. To "undo" multiplication, we divide, so 3y/3 is just y; 4x/3 is 4/3 x; and 12/3 is 4. So the new equation when we solve for y is: y = 4/3 x + 4. Is there any part of that you don't follow?

OpenStudy (anonymous):

no iam understanding it all thank you

OpenStudy (imstuck):

Great! Next thing is for you to decide what the slope of that line is. Do you know it?

OpenStudy (anonymous):

no

OpenStudy (imstuck):

Ok the y is just the y; the number in front of the x (which is the "m" in y = mx + b) is the slope, and the "b" is the y- intercept. So if the number in front of the x is the slope, what is that number in the equation we just solved for y?

OpenStudy (imstuck):

y = 4/3 x + 12 is the equation

OpenStudy (anonymous):

are you saying that is the answer

OpenStudy (imstuck):

No, I'm saying that is the slope of the equation we just solved for y, which was the given equation. The "m" is the slope. What is the slope?

OpenStudy (imstuck):

Look at it like this, maybe it will help some: y = m x + b y = 4/3 x + 12

OpenStudy (anonymous):

so to get m i divide 4 and 3

OpenStudy (imstuck):

The m is the slope of the equation. In this case, the slope is 4/3 or to type it in words, four thirds. Four thirds is the slope of the given line.

OpenStudy (imstuck):

You can divide it out and get 1.3333333333333 forever, but you don't need to do that. You can just leave it in fraction form. It doesn't make it harder to leave it in fraction form. It won't matter at all.

OpenStudy (anonymous):

Would the answer be y=4/3x+3

OpenStudy (imstuck):

You're close, but remember that the equation is found from the point-slope formula which is \[y -y _{1}=m(x-x _{1})\]

OpenStudy (imstuck):

The y doesn't change from a y, it's the \[y _{1}\] that changes according to what your points are. Here, your y point is 4, right? And you x point is -3. So use those and the slope you just found and fill them into the equation. What do you get?

OpenStudy (imstuck):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (imstuck):

\[y-4=\frac{ 4 }{ 3 }(x-(-3))\]

OpenStudy (imstuck):

\[y-4=\frac{ 4 }{ 3 }(x+3)\]

OpenStudy (imstuck):

Now you have to distribute the 4/3 through the parantheses. You get:\[y-4=\frac{ 4 }{ 3 }x+4\]

OpenStudy (imstuck):

To get the y alone, you add 4 to both sides:\[y=\frac{ 4 }{ 3 }x+8\]

OpenStudy (anonymous):

you get y=4/3x+8

OpenStudy (imstuck):

YES! see how smart you are and didn't even know it!

OpenStudy (imstuck):

That equation is the answer. It has the same slope of 4/3 and goes through your given points. Do you understand all that?

OpenStudy (anonymous):

Thank you so much for your help.

OpenStudy (imstuck):

You are so welcome! I love to help with math!

OpenStudy (imstuck):

Anytime!

OpenStudy (anonymous):

Thank you